How many fissions take place per second in a 190-MW reactor? Assume 200 MeV is released per fission.
Answer #1

First let's see how many Mev we need to equal 190 megawatts, or 1.9 * 10^8


watts, or joules per second. One watt equals 1 joule (1 volt-coulomb) per sec.


Since 1 coulomb = 6.28 * 10^18 electrons, 1 joule = 6.28 * 10^18 eV.


So 190 megawatts = (1.9 * 10^8)(6.28 * 10^18) = 11.932 * 10^26 eV per sec.


If each fission produces 200 MeV or 2.0 * 10^8 eV, we need:


(11.932 * 10^26) / (2.0 * 10^8) = 5.966 * 10^18 fissions per second.